OpenStudy (dtan5457):

Let f(n) be the function that assigns to each integer greater than 1 the sum of its distinct prime factors. For example, since 2 and 5 are the only prime factors of 20, f(20)=7, and since 5 is the only prime factor of 25, f(25)=5. Which of the following statements are true? f(2n)=f(n) f(2n)=f(n)+2 f(2n)=2f(n)

OpenStudy (empty):

for the first one let n=3. for the second one let n=2. for the third one let n = 15.

OpenStudy (danica518):

f(2n)=f(n)+2

OpenStudy (danica518):

i am new to os, are we allowed to give out answers

OpenStudy (empty):

Right, so if n=2 then we have: f(4)=f(2)+2 which evaluates to 2=2+2 So true or false?

ganeshie8 (ganeshie8):

f(ab) = f(a)+f(b) whenever gcd(a,b)=1

ganeshie8 (ganeshie8):

so we may say f(2n) = f(2) + f(n) whenever n is odd

ganeshie8 (ganeshie8):

wait a second, i think i have interpreted the question incorrectly it is about "adding the distinct prime factors", not just number of distinct primes

ganeshie8 (ganeshie8):

I think it shouldn't matter either way since both are logarithm alike !